# Generalized Animal Imitator: Agile Locomotion with Versatile Motion Prior

### v2

@@ III-C Imitation Reward for Functionality and Style
Stylization Reward Scheduling: To learn the style quickly and stably, we propose to use both adversarial stylization reward and joint angle tracking reward with a balanced scheduling mechanism. Considering the discriminator as a "coach", we utilize the mean adversarial reward as an indication of how the coach is satisfied with the current performance. When it’s not satisfied with the current performance of the robot, it provides detailed instructions for the robot to learn. Specifically, our stylization reward follows:
$r_{\text{style}}(s_{t},\hat{s}_{t}^{\text{ref}})$ $=w_{{\text{style}}}^{\text{adv}}*r_{\text{style}}^{\text{adv}}+w_{\text{style}}^{\text{joint}}*r_{\text{style}}^{\text{joint}}$
$+w_{{\text{style}}}^{\text{adv}}*(1-\displaystyle\mathop{\mathbb{E}}_{s_{t}\in S}(r_{\text{style}}^{\text{adv}}(s_{t},s_{t+1})))*r_{\text{style}}^{\text{joint}}$
With the given formulation, our stylization reward provided dense rewards during the beginning of training, enabling the robot to quickly catch the essence of different agile locomotion skills. Our stylization reward also provides detailed instruction as the training proceeds, enabling the robot to refine its gait and lead to more stable training.
III-D Solving Downstream Tasks with Motion Prior:
@@
Figure 4: Solving High-level Tasks with Our Motion Prior. Our high-level policy takes the high-level task information and the proprioception as input and output high-level latent command for low-level policy.
@@
For intricate tasks like jumping over gaps, addressing them from scratch presents challenges including acquiring necessary agile locomotion skills, such as jumping and rapid running within limited interactions, and the intensive engineering needed to harmonize the reward for top-tier tasks while regularizing the robot’s motion. With a low-level motion prior, robots can instantly harness existing skills encapsulated within the prior and channel their efforts into high-level strategizing. For each distinct downstream task, we train a high-level policy $\pi_{\text{high}}$ (As shown in Figure 4) takes the high-level observation $\mathbf{o}_{\text{high}}$, and proprioception of the robot and outputs latent command for low-level motion prior to utilize the existing agile locomotion skills: $a_{t}=\pi_{\text{low}}(\pi_{\text{high}(\mathbf{o}_{\text{high}},s_{t}),\mathbf{E}_{\text{prop}}(s_{t})})$.
III-E Implementation Details:
Observation Space: Our low-level observation includes joint angles, joint velocities, gravity vector in the robot frame, and the previously executed actions.
Action Space: Our action space is defined as target joint angles for 12 joints of our robot. The target joint angles are converted to torque command with PD controller where KP=40, KD=1.0. Our control frequency in simulation and real-world is 25Hz.
Reference Encoder $E_{ref}$: Our reference encoder is a two-layer MLP with $[256]$ hidden units, mapping the reference motion segment into a $64$ dimensional latent distribution.
Proprioception Encoder $E_{prop}$: Our reference encoder is a two-layer MLP with $[256]$ hidden units, mapping proprioception information into a $64$-dim robot state feature.
Low-level Policy $\pi_{low}$ and Value Network $V_{low}$: Our low-level policy is a three-layer MLP with $[256,128]$ hidden units, mapping the robot state feature and latent command to $12$-dim robot target joint angles. Our low-level value network shares the same structure with an additional $64$-dim motion embedding as input, and output 1-dim value for RL training.
High-level Policy $\pi_{high}$ and Value Network $V_{high}$: Our high-level is formulated as a three-layer MLP with $[256,128]$ hidden units, mapping proprioception information and high-level task information to high-level latent command for low-level motion prior. Our high-level value network shares the same structure. High-level task information depends on specific task.
Low-level training termination: As stated in Sec 3, we terminate the episode when the position error between the current robot position and root position of reference motion is larger than $1(^{2})$, or the pose error between current robot pose and root pose of reference motion is larger than $\frac{1}{2}\pi$. We set the maximum simulation time for each episode to $20$ which corresponds to maximum episode length = 500.
Simulation Setup: We utilize IsaacGym[32] to simulate 4096 robots in parallel. During motion prior training, for each robot, we uniformly sample a reference motion from the dataset for it to track.
IV Experiments
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
@@ IV Experiments
Figure 6: Real World Backflip Trajectory Comparison: Each row represents a single trajectory (From top to bottom: Reference Motion, VIM, GAIL, Motion Imitation). Trajectories are shown from left to right.
@@
We evaluate our system in simulation and real-world, benchmarking it against prior work for low-level locomotion skill learning and various High-level tasks. Our robot demonstrates life-like agility in the real world.
@@
TABLE II: Evaluation of Motion Prior in Simulation: We compare Horizontal and Vertical Root Position (Root Pos (XY), Root Pos (Height)), Root Orientation (Root Ori), Joint Angle, and End Effector Position (EE Pos) tracking errors and RL objectives of all methods. Our methods outperform all baselines in terms of smaller tracking errors, higher episodic returns, and longer episode lengths. GAIL baseline shows a smaller root position tracking error since it can’t follow the reference motion leading to early termination of the episode.
@@
TABLE III: Evaluation of Motion Prior in Real (Left): We collect representative metrics for different locomotion skills with corresponding metrics from reference motion. $N/A$ denotes completely failed skills in real.
@@
Figure 7: Latent Skill Space t-SNE. We visualize the latent skill space by visualizing the latent embedding for varying motion segments.
Figure 8: High-level Tasks Evaluation in Simulation: Solid line and shaded area denote the mean and std across random seeds. Our system outperforms all baselines on all tasks.
IV-A Evaluation of Learned Motion Priors
Our system’s proficiency in learning a range of agile locomotion skills from the reference motion dataset (discussed in Sec III-A) is initially assessed.
Baselines: We benchmark our method against two representative baselines: Motion Imitation [40, 5, 22] baseline represents a thread of recent works whose imitation rewards are defined solely with errors between current robot states and the corresponding reference states. Generative Adversarial Imitation Learning (GAIL) baseline represents a thread of recent work [42], whose imitation reward is solely provided by the discriminator trained to distinguish trajectories generated by the policy from the ground truth reference motions. WASABI baseline represents a modified version of WASABI [30] for our setting. Given that our reference motions consist only of state sequences, they offer less supervision compared to expert action sequences, rendering motion prior learning more challenging. Each method trains for $2\times 10^{9}$ iterations across 3 random seeds. Both our method and the Motion Imitation baseline adopt identical reward scales for all motion error-tracking rewards. Likewise, our approach and GAIL maintain the same scale for adversarial stylization rewards.
TABLE IV: High-level Tasks in Real World: We compare Following Command + Jump Forward policies of all methods in real. $N/A$ denotes completely failed skills in real. Our methods outperform all baselines in real for most metrics.
@@ IV-A Evaluation of Learned Motion Priors
Simulation Evaluation: In the simulation, we measure average imitation tracking errors for various agile locomotion skills, episode returns, and trajectory lengths across random seeds. Specifically as listed in Table II, where the tracking error of root pose represents the ability of the robot to reproduce the locomotion skill, and the tracking error of joint angle and end effector position represents the ability of the robot to mimic the style of reference motion. Our method achieves a similar root pose tracking error as the motion imitation baseline with a much smaller joint angle tracking error. This shows our method striking a balance between functionality and style, superior to the motion imitation baseline that focuses solely on functionality. Meanwhile, the GAIL baseline failed to learn the functionality of the reference motions which leads to short episode length and the least episode return. We surmise that the GAIL baseline’s inadequacy arises for two main reasons: First, exclusive reliance on adversarial stylization reward does not offer temporally consistent guidance throughout skill learning due to misaligned rewards across timesteps. Second, the mode collapse issue inherent in adversarial training hinders the robot from mastering highly agile skills, such as backflipping. The shortcomings of the Motion Imitation baseline may stem from the challenges of balancing different terms and selecting suitable hyperparameters when concurrently learning multiple agile locomotion skills. Comparing our VIM with and without stylization reward scheduling, we find the former exhibits enhanced style tracking performance, underscoring the value of stylization reward scheduling in refining robot gait tracking.
@@
Real World Evaluation: We gauge learned agile locomotion skills in real-world scenarios. Due to the lack of precise robot pose, we resort to specific metrics tailored to different locomotion skills, detailed in Table III. For Jump While Running & Jump Forward & Jump Forward (Syn) & Backflip, we measure the jumping height and jumping distance. For Pace & Canter & Walk & Trot and Left Turn & Right Turn, we measure the linear and angular velocity, respectively. Results reveal that our method retains most of the reference motion functionality. The only significant deviation, observed in the Canter motion, arises from inherent differences between animal movement (its source) and our robot’s capabilities since quadrupeds could utilize tendons at their ankle to achieve higher running speed with the specified motion which is not possible for our robot. Even with comparable root pose tracking errors in simulations, our method outshines the Motion Imitation baseline in real-world metrics like jumping height, distance, and velocity tracking error. This suggests that mirroring the style of the reference motion improves sim2real transfer for natural gaits. GAIL baseline struggled to reproduce most real-world locomotion skills. A visual comparison of real-world trajectories is available in Figure  5 and Figure 6, showing our method’s superiority in capturing both motion functionality and style. Due to the poor performance of WASABI baseline in simulation, we didn’t evaluate it in the real world.
@@
Latent Skill Space Visualization: To better understand our low-level motion prior, we visualize the learned latent skill space in Figure 8 by visualizing the latent embedding corresponding to motion segments in our reference motion dataset via t-SNE [53]. We find that different skills are separated into different regions with clear boundaries. Our reference motion encoder also clusters the skills with similar semantic meaning together: the embeddings from Left Turn sequence are close to the embeddings from Right Turn sequence, which enables the smooth transition between different skills, the embeddings from Jump While Running & Jump Forward & Jump Forward (Syn) sequence are clustered together. These observations suggest that our system learned a smooth and semantically meaningful latent skill space for solving High-level tasks.
TABLE V: Ablation Study of Motion Prior in Simulation: We compare our method and two variant of our method without Functionality (Func) reward or Style reward on all different objectives

### v3

@@ 3.2 Imitation Reward for Functionality and Style
Stylization Reward Scheduling: To learn the style quickly and stably, we propose to use both adversarial stylization reward and joint angle tracking reward with a balanced scheduling mechanism. Considering the discriminator as a "coach", We utilize the mean adversarial reward as an indication of how the coach is satisfied with the current performance. When it’s not satisfied with the current performance of the robot, it provides detailed instructions for the robot to learn. Specifically, our stylization reward follows: $r_{\text{style}}(s_{t},\hat{s}_{t}^{\text{ref}})=w_{{\text{style}}}^{\text{adv}}*r_{\text{style}}^{\text{adv}}+w_{\text{style}}^{\text{joint}}*r_{\text{style}}^{\text{joint}}+w_{{\text{style}}}^{\text{adv}}*(1-\displaystyle\mathop{\mathbb{E}}_{s_{t}\in S}(r_{\text{style}}^{\text{adv}}(s_{t},s_{t+1})))*r_{\text{style}}^{\text{joint}}$ With this formulation, our stylization reward provides dense rewards at the beginning of training, enabling the robot to quickly catch the essence of different agile locomotion skills, and provides detailed instruction as the training proceeds, enabling the robot to refine its gait.
3.3 Solving Downstream Tasks with Motion Prior:
@@
Figure 4: Solving High-level Tasks with Our Motion Prior. Our high-level policy outputs high-level latent command for low-level policy.
@@
For intricate tasks like jumping over gaps, starting from scratch is challenging due to the need for agile locomotion skills and the intensive engineering to balance rewards and regularize motion. Using a low-level motion prior, robots can immediately use existing skills and focus on high-level strategies. For each distinct downstream task, we train a high-level policy $\pi_{\text{high}}$ (As shown in Figure 4) takes the high-level observation $\mathbf{o}_{\text{high}}$, and proprioception of the robot and outputs latent command for low-level motion prior to utilize the existing skills: $a_{t}=\pi_{\text{low}}(\pi_{\text{high}(\mathbf{o}_{\text{high}},s_{t}),\mathbf{E}_{\text{prop}}(s_{t})})$.
Additional implementation details about observation/action space, reference/proprioception encoder, low-level/high-level policy, and value network can be found in the Appendix F.
4 Experiments
@@
Figure 5: Real World Backflip Trajectory: Each row represents a single trajectory (From top to bottom: Reference Motion, VIM, GAIL, Motion Imitation). Trajectories are shown from left to right.
@@
We evaluate our system in simulation and real-world, comparing with prior work for low-level skill learning and various high-level tasks. Our robot demonstrates life-like agility in the real world.
@@
Table 2: Evaluation of Motion Prior in Simulation: We compare Horizontal and Vertical Root Position (Root Pos (XY), Root Pos (Height)), Root Orientation (Root Ori), Joint Angle, and End Effector Position (EE Pos) tracking errors and RL objectives of all methods. Our methods outperform all baselines in terms of smaller tracking errors, higher episodic returns, and longer episode lengths. GAIL baseline shows a smaller root position tracking error since it can’t follow the reference motion leading to early termination of the episode.
@@
VIM (w/o Style Reward) $1.49\scriptstyle{\pm 0.69}$ $0.00\scriptstyle{\pm 0.01}$ $0.12\scriptstyle{\pm 0.06}$ $5.14\scriptstyle{\pm 1.71}$ $0.25\scriptstyle{\pm 0.06}$ $6.28\scriptstyle{\pm 6.73}$ $109.76\scriptstyle{\pm 110.67}$
Table 3: Evaluation of Motion Prior in Real: We collect representative metrics for different skills with corresponding metrics from reference motion. $N/A$ denotes completely failed skills in real.
@@
Figure 6: Latent Skill Space t-SNE. We visualize the latent embedding for varying motion segments.
Figure 7: High-level Tasks Evaluation in Simulation: Solid line and shaded area denote the mean and std across random seeds. Our system outperforms all baselines.
4.1 Evaluation of Learned Low-level Motion Priors
Baselines: We benchmark our method against three representative baselines: Motion Imitation [4, 17, 20] baseline represents a thread of recent works whose imitation rewards are defined solely with errors between current robot states and the corresponding reference states. Generative Adversarial Imitation Learning (GAIL) baseline represents a thread of recent work [18], whose imitation reward is solely provided by the discriminator trained to distinguish trajectories generated by the policy from the ground truth reference motions. WASABI baseline represents a modified version of WASABI [23] for our setting. Each method trains for $2\times 10^{9}$ samples across 3 random seeds. Both our method and the Motion Imitation baseline adopt identical reward scales for all motion error-tracking rewards.
Table 4: High-level Tasks in Real World: We compare Following Command + Jump Forward policies of all methods in real. $N/A$ denotes completely failed skills in real. Our methods outperform all baselines in real for most metrics.
@@ 4.1 Evaluation of Learned Low-level Motion Priors
Simulation Evaluation: In the simulation, we measure average imitation tracking errors, episode returns, and trajectory lengths across random seeds. As listed in Table 2, the tracking error of root pose represents the ability of the robot to reproduce the locomotion skill, and the tracking error of joint angle and end effector position represents the ability of the robot to mimic the style of reference motion. Our method achieves a similar root pose tracking error as the motion imitation baseline with a much smaller joint angle tracking error. This shows that our method strikes a balance between functionality and style, superior to the motion imitation baseline that focuses mainly on functionality. Meanwhile, the GAIL baseline failed to learn the functionality of the reference motions leading to short episode length and the least episode return. We surmise that the GAIL baseline’s inadequacy arises from the adversarial reward does not offer temporally consistent guidance throughout skill learning and the mode collapse issue inherent in adversarial training hinders the robot from mastering highly agile skills, such as backflipping. The poor performance of the Motion Imitation baseline may stem from the challenges of balancing different terms and selecting suitable hyperparameters when concurrently learning multiple agile locomotion skills.
Ablation Study of Learned Motion Prior: We provide the ablation study over the reward term and the scheduling mechanism as shown in Table 2. We found that without Functionality reward, the learned controller could not robustly track the reference motion resulting in smaller Episode Return and, shorter Episode Length on the other hand, removing Style reward results in a significantly higher Joint Angle and End-Effector tracking error. Comparing VIM with and without stylization reward scheduling, we find the former exhibits enhanced style tracking performance, underscoring the value of stylization reward scheduling in refining robot gait tracking.
@@
Real World Evaluation: We evaluated learned agile locomotion skills in the real world using specific metrics tailored to different skills, as detailed in Table 3. We repeated our experiment three times per skill per method per seed since our real-world experiment. For Jump While Running/Jump Forward/Jump Forward (Syn)/Backflip, we measured jumping height and distance. For Pace/Canter/Walk/Trot and Left Turn/Right Turn, we measured linear and angular velocity. Results show our method retains most of the reference motion functionality. The only significant deviation observed in Canter is due to differences between animal and robot capabilities, as quadrupeds use tendons to achieve higher running speeds, which our robot lacks. Despite similar root pose tracking errors in simulations, our method outperforms the Motion Imitation baseline in real-world metrics like jumping height, distance, and velocity tracking error, indicating that mirroring reference motion style improves sim2real transfer for natural gaits. The GAIL baseline struggled with real-world locomotion skills. Figure 5 visually compares real-world trajectories, showing our method’s superiority in capturing both motion functionality and style. Due to poor simulation performance, the WASABI baseline was not evaluated in the real world.
@@
Latent Skill Space Visualization: We visualize the learned latent skill space in Figure 7 by visualizing the latent embedding corresponding to motion segments in our reference motion dataset via t-SNE [58]. We find that different skills are separated into different regions with clear boundaries. Our reference motion encoder also clusters the skills with similar semantic meaning together: embeddings from Left Turn / Right Turn sequence are close, which enables the smooth transition between different skills. embeddings from Jump While Running & Jump Forward & Jump Forward (Syn) sequence are clustered together. These observations suggest that our system learned a smooth and semantically meaningful latent skill space for solving high-level tasks.
4.2 Evaluation on High-level Tasks
To evaluate how our method leverages learned agile locomotion skills for high-level tasks, we designed a set of tasks and tested our method against baselines in simulation and the real world.
High-level Tasks & Observation: Our tasks include: Following Command: directing the robot to move with specific linear and angular velocities. Linear velocity commands range from $0\sim 2$ m/s, and angular velocity commands range from $-2\sim 2$ rads/s. In our motion prior, the robot is trained to move and turn at the reference motion’s speed. Hence, to follow a command precisely, the high-level policy needs to smoothly interpolate between different speeds. Jump Forward: directing the robot to jump while running. We have adapted a subset of jumping rewards from CAJun [59] to evaluate policy interpolation between jumping and running motions within a fixed timeframe. Following Command + Jump Forward: directing the robot to either jump forward or adjust to changing commanded speeds. To optimize episode return, the robot should not only use the agile locomotion skills from the reference motion dataset but also develop unobserved skills like executing sharp turns. Detailed high-level observations for different tasks are provided in the Appendix F.
Baselines: Given the baseline’s subpar performance in low-level motion prior training, we compare our system with three representative baselines without pre-trained low-level controllers: PPO [54]: Controllers trained exclusively on high-level task rewards. AMP [22]: Utilizes reference motion for styling reward in adversarial imitation learning and learns high-level tasks while mimicking reference motions. Hierarchical Reinforcement Learning (HRL) from Jain et al. [60]: Learns a high-level policy sending latent commands to a low-level controller, resembling works that decompose tasks into sub-problems [61, 62, 63, 64, 65, 66, 67]. For fair comparison, we removed the trajectory generator in [60], used PPO for AMP and HRL, and used full reference motion for AMP and HRL with AMP.
Evaluation in Simulation & Real World: We trained all methods on each high-level task for $4\times 10^{8}$ samples with 3 random seeds. Simulation results are detailed in Figure 7, and real-world results are provided in Table 4. Real-world Following Commands trajectory is also provided in Appendix I. For the Following Command task, all methods mastered basic locomotion, but ours excelled in efficiency and smooth transitions between diverse linear and angular velocities. In the Jump Forward and Following Command + Jump Forward tasks, which required advanced jumping abilities, baselines struggled. They either moved forward continuously, remained grounded when prompted to jump, or toppled to avoid energy consumption penalties. In contrast, our system seamlessly integrated jumping and running actions, achieving the highest episode return. Despite having a comprehensive reference motion dataset, baselines couldn’t harness the skills effectively. This likely stems from the difficulty of deriving agile locomotion skills using only adversarial stylization rewards, similar to the GAIL baseline’s poor performance in low-level training.
5 Limitations
Our current system exhibits several limitations: 1) Safety is not assured with our current method. Introducing safety constraints during training could mitigate this issue; 2) The robot’s limited capacity restricts its ability to fully replicate certain motion capture data, like cantering. Upgrading the hardware could address this limitation; 3) Currently, our system does not incorporate dynamics information. This could be improved by integrating adaptation techniques during deployment; 4) The system’s low-level motion priors and high-level policies currently lack perceptual information.
6 Conclusion
In this paper, we propose Versatile Instructable Motion prior (VIM) which learns agile locomotion skills from diverse reference motions with a single motion prior. Our simulation and real-world results show that our VIM captures both the functionality and style of locomotion skills from reference motions. Our VIM also provides a temporally consistent and compact latent skill space representing different locomotion skills for high-level tasks. With agile locomotion skills in our VIM, complex High-level tasks can be solved efficiently with minimum human effort.
Acknowledgement
This work was supported, in part, by NSF CCF-2112665 (TILOS), NSF 1730158 CI-New: Cognitive Hardware and Software Ecosystem Community Infrastructure (CHASE-CI), NSF ACI-1541349 CC*DNI Pacific Research Platform.
Appendix
Appendix A Reference Motion Dataset
Our reference motions (11 reference motions in total) come from motion capture of animal motion, trajectory optimization method, and synthesized data with a generative model. The length of our reference motion ranges from 32 to 500. During training, we repeat the reference motions cyclically to fit the length of the episode.
Table 5: Reference Motion Dataset:
Motion length 500 32 500 38 45 38
Motion length 500 500 129 120 64
Appendix B Performance Across Different Reference Motions
In our framework, when the root pose of the simulated robot diverges too much from the reference root pose, we terminate the episode, as described in Sec 3 . In this case, the episode length is a good indicator of whether the learned policy could follow the reference motions. As shown in Fig. 8 , the performance of our framework varies when imitating different reference motions. When imitating relatively steady motions like Walk (Mocap), Pace (Mocap), Left Turn (Mocap), the learned controller could track the motion for a longer period. When imitating relatively agile motions, especially with high moving speed, such as Canter (Mocap), Jump While Running (Mocap), the performance of our system drops. This phenomenon is rooted in the methodology disparities between our robot and real animals.
Appendix C Additional Discussion about ASE
We would like to clarify the significant differences between our method and the ASE baseline in the following aspects:
  • Skill Control and Learning: Unlike ASE, which learns motor skills from a transition dataset in an unsupervised manner without controlling the outcome, our method intentionally learns specific motor skills from our reference motion dataset. This controlled learning approach ensures that critical skills, such as the jump motion are effectively acquired. This capability is crucial for constructing a motion prior tailored for varied high-level tasks. As shown in ASE video and demonstration even though there is jump motion in their dataset, ASE failed to learn it.
  • Long-Term Skill Acquisition: ASE focuses on learning motor skills at the transition level, limiting its ability to learn complex, long-term motor skills like backflipping, which require extended motion sequences. Our method, however, leverages a combination of adversarial styling and dense tracking rewards, providing structured supervision for acquiring long-term motor skills, including agile locomotion abilities like backflipping and jumping forward. This sequence-level modeling is essential for the effective learning of complex locomotion skills.
  • Performance Evaluation: Since ASE learns different motor skills in an unsupervised manner, it’s difficult to evaluate the performance of the learned low-level controller (There is no evaluation or benchmark for the low-level controller in ASE paper). While our method learns to imitate different locomotion skills in the dataset at the sequence level, we could directly benchmark the tracking error to evaluate the quality of the learned low-level controller. Benchmark over the low-level controller is also important if we want to build a backbone of low-level skills for diverse potential high-level tasks.
Appendix D Additional Discussion over Skill Learning Frameworks
In this section, we provide further discussion on the existing skill-learning framework:
  • Function Tracking: The resulting controller of a skill-learning framework can accurately track the movement of the robot’s base.
  • Skill Tracking: The resulting controller can faithfully replicate the joint movement patterns of the robot.
  • Agility: The controller is capable of producing highly agile locomotion skills. Since there is no universally accepted definition of "agile", in our work, we consider a skill "agile" if it involves the robot leaving the ground, such as in a backflip or jump, or running/turning at high speed.
  • Control Skills to Learn: Given a fixed set of reference motions, the resulting controller can reliably reproduce specific skills. Unsupervised methods like ASE do not guarantee performance on any particular skill in the dataset.
  • Multiple Skills: The resulting controller is capable of performing a variety of different skills.
  • Diverse Sources: The resulting controller learns different skills from various sources.
  • Reusable: The resulting controller can be reused for tasks beyond reproducing the reference motion’s skills.
  • No Privileged Information: The resulting controller does not require privileged information (such as the robot’s velocity or position in the world frame) during deployment.
  • Real Deployment: The proposed framework is validated in real-world scenarios.
Additional discussion on the performance of ASE/AMP-based methods in agile locomotion skills: ASE struggles to capture agile locomotion skills for the following reasons:
  • Difficulty in Learning Agile Skills: Agile locomotion skills, such as jumping and backflipping, are inherently more challenging to learn compared to other locomotion skills like walking or trotting. Due to the well-known mode-collapse issue in the generative adversarial learning paradigm, it is particularly difficult for generative adversarial methods (like ASE/AMP) to discover and learn these complex skills in an unsupervised manner.
  • Limitations of Transition-Level Learning: As discussed in Appendix C, ASE/AMP performs adversarial learning at the transition level, focusing on the current and previous states of the robot (as shown in Formula (3) in ASE[18]). However, skills that require a longer sequence of actions, such as backflipping or jumping, are difficult to learn with this approach. For example, executing a backflip involves multiple stages: sitting down, lifting the front legs, pushing off with the rear legs, adjusting the pose in mid-air, and landing. Similarly, jumping and running require coordinated stages of movement. Transition-level supervision lacks the long-term guidance needed to learn these complex, agile skills.
  • Limited Input for Discriminator: ASE/AMP discriminators only consider joint angles as input, making it more challenging for the robot to learn agile locomotion skills that involve significant changes in the robot’s position and orientation.
Appendix E Single Skill Comparison in Simulation
We conducted additional evaluations focusing on single skill learning, where all methods are required to learn a single reference motion using identical hyperparameters. For these experiments, we removed the motion embedding from the critic since there is only one skill to learn. We selected Jump Forward (Optimization), and Jump while Running (Mocap) as representative skills for agile locomotion, and Trot (Mocap) and Left Turn (Mocap) as representative skills for normal locomotion. Each method was trained with $2\times 10^{9}$ samples per skill across three random seeds.
In general, as shown in Table 6, single-skill tracking tends to deliver better results, in terms of longer episode length and higher episode return (representing the overall performance), for the specified skill because the task is easier to learn and more samples are dedicated to that particular skill. (In our low-level motion prior training stage, all skills share the total number of samples.) For agile locomotion skills, our method outperforms both GAIL (Single Skill AMP) in terms of better tracking of the robot’s root movement and the motion imitation baseline by achieving smaller joint tracking errors. For normal locomotion skills, all methods deliver reasonable results for both Trot (Mocap) and Left Turn (Mocap). However, GAIL (Single Skill AMP) exhibits slightly higher joint tracking error for Trot, which we attribute to the lack of temporal alignment in the adversarial reward. Although the GAIL baseline can faithfully reproduce the skill, it shows a slightly higher tracking error. It’s important to note that the results in Table 6 should not be directly compared with those in Table 2 as longer episodes tend to accumulate more errors, leading to larger tracking errors, and the experiment setting is not identical.
Table 6: Evaluation of Motion Prior in Simulation for single Reference motion: We compare Horizontal and Vertical Root Position (Root Pos (XY), Root Pos (Height)), Root Orientation (Root Ori), Joint Angle, and End Effector Position (EE Pos) tracking errors and RL objectives of all methods.